The P/E ratio for Mailbu Boats (MBUU) is 13.24 and the expected growth rate is 16%. If the plowback ratio is 80% what is the current required rate of return (market capitalization rate) of the stock?
Answer #1

P/E Ratio = 13.24

Price = 13.24E

As per Constant Growth Model,

Price = 0.20(E)(1.16)/(r - 0.16)

r = 0.20(E)(1.16)/(13.24)(E) + 0.16

r = 17.75%

Required Rate = 17.75%